OpenStudy (anonymous):

Alicia can row 5 miles downstream in the same time it takes her to row 2 miles upstream. She rows downstream 4 miles/hour faster than she rows upstream. Find Alicia's rowing rate each way. Round your answers to the nearest tenth, if necessary.

OpenStudy (anonymous):

5/x = 2/x +4 5/x = (2+4x)/x 5=2+4x x=.75 downstream rate =5/.75=6.7miles per hr upstream rate= 2/.75=2.7miles per hr